OpenStudy (anonymous):

The perimeter of a rectangle is 228 cm the length is 39 less than twice the width find dimension.

OpenStudy (anonymous):

ui pamea nasaagut na sakin yan...

OpenStudy (anonymous):

need second opinion :))

OpenStudy (anonymous):

\[l=2w-39\] \[2l+2w=228\] \[2l+l+39=228\] \[3l=189\] \[l=63\]

OpenStudy (anonymous):

nevermind i know the answer already ^_^

OpenStudy (anonymous):

thanks by the way ^_^

OpenStudy (anonymous):

first equation: 2x + 2y = 228, where x=width and y=length... second equation y=2x-39...solve simultaneously and get that x=51 and y=63

OpenStudy (anonymous):

thank you ^_^